OpenStudy (anonymous):

In the December 9, 2008, New York Times article, "Fedex lowers 2009 Earnings Forecast," Fedex said it expeceted to earn an average of $4.13 a share for the fiscal year ending in May, down from its original estimate of an average of $5.00 a share. Analysts were predicting that the company would earn $5.15 a share according to Thomas Reuters. What percent reduction is the company's new estimate from the original estimate in earnings per share? What percent reducation in earnings per share is there between the orginal estimate and the analysts perdictions? Round to the nearest tenth percent.

OpenStudy (rulnick):

change = (new - old) / old = (4.13-5) / 5 = -17.4%

OpenStudy (rulnick):

Second question is worded a little funky. Taken literally, it would be a "negative reduction," a "reduction of -3.0%"

OpenStudy (rulnick):

Do you think it meant to say "from the analysts predictions to the original estimate"? If so, it would be a reduction of (5.15-5)/5.15 = 2.9%.